Dell Corporation makes two types of tablets, A and B in equal quantities. That is, for every unit of tablet A that firm produces, it also produces 1 unit of tablet B. Dell corporation total cost function is
TC = 100 + Q + 2Q
Where Q is the number of units of output(Where each unit contains 1 unit of tablet A and 1 unit of tablet B), and TC is the total cost (in$). The demand curves for the firm two products are:
Pa = 200-Qa
Pb = 150 - 2Qb
Where Pa and Qb are the price and output A, and Pb and Qb are the price and output of product B.
a. How much of each product should the Dell Corporation produce and sell per period of time?
b. What price should it charge for each product ?
1
Expert's answer
2018-08-03T15:42:08-0400
TC = 100 + Q + Q^2
Pa = 200-Qa
Pb = 150 - 2Qb
Qa = Qb, Qa + Qb = Q,
a. The Dell Corporation should produce and sell per period of time such quantity, for which MR = MC.
For product A: MR = TR' = 200 - 2Qa, MC = TC' = 1 + 2Q = 1 + 4Qa, so 200 - 2Qa = 1 + 4Qa,
6Qa = 199,
Qa = Qb = 33 units.
b. The price it should charge for each product is:
Pa = 200 - Qa = 200 - 33 = $167,
Pb = 150 - 2Qb = 150 - 2*33 = $84.